If F(4)=0 and F(6)=6, which of the following could represent F(x)?
I know the answer is 3x-12 but don't know how they came to that from what I am given.

Answer by stanbon(75887)

If F(4)=0 and F(6)=6, which of the following could represent F(x)?
I know the answer is 3x-12 but don't know how they came to that from what I am given.
--------------
You have two points: (4,0) and (6,6)
The form of the linear equation thru these point is y=mx+b
m = (6-0)/(6-4) = 6/2 = 3
So, you have m=3 and y = 6 when x=6.
6 = 3*6 + b
b = -12
----------
EQUATION:
y = 3x-12
=============
Cheers,
Stan H.